# Zero-Shot Trajectory Planning for Signal Temporal Logic Tasks

### v1

@@ 3.2 Decompositions of STL Formulae
  • Invariance Progress: we denoted by $\mathcal{I}(a_{\Lambda},b_{\Lambda},\mu)$ that $\forall t\in[a_{\Lambda},b_{\Lambda}],\mathbf{x}_{t}\vDash\mu$.
Note that we use subscript ${\Lambda}$ in time interval $[a_{\Lambda},b_{\Lambda}]$ as $a_{\Lambda}$ and $b_{\Lambda}$ may not be a fixed time and involve variables in $\Lambda$ subject to time constraints. Therefore, the STL formula $\varphi$ is decomposed into a tuple $(\mathbb{P}_{\varphi},\mathbb{T}_{\varphi})$, where $\mathbb{P}_{\varphi}$ is the set of progresses and $\mathbb{T}_{\varphi}$ is the set of time constraints over variable set $\Lambda$. Such decomposition is defined recursively as follows:
  • If $\varphi=\mathrm{F}_{[a,b]}\mu$, then we have $\mathbb{P}_{\varphi}=\{\mathcal{R}(\lambda_{i},\lambda_{i},\mu)\}$ and $\mathbb{T}_{\varphi}=\{\lambda_{i}\in[a,b]\}$, where $\lambda_{i}$ is a new time variable.
  • If $\varphi=\mathrm{G}_{[a,b]}\mu$, then we have $\mathbb{P}_{\varphi}\!=\!\{\mathcal{I}(a,b,\mu)\}$ and $\mathbb{T}_{\varphi}\!=\!\emptyset$.
  • If $\varphi=\mu_{1}\mathrm{U}_{[a,b]}\mu_{2}$, then we have $\mathbb{P}_{\varphi}\!=\!\{\mathcal{I}(a,\lambda_{i},\mu_{1}),$ $\mathcal{R}(\lambda_{i},\lambda_{i},\mu_{2})\}$ and $\mathbb{T}_{\varphi}=\{\lambda_{i}\in[a,b]\}$, where $\lambda_{i}$ is a new time variable.
  • If $\varphi^{\prime}=\varphi_{1}\wedge\varphi_{2}$, then we merge the progresses and time constraints, i.e., $\mathbb{P}_{\varphi}=\mathbb{P}_{\varphi_{1}}\cup\mathbb{P}_{\varphi_{2}}$ and $\mathbb{T}_{\varphi}=\mathbb{T}_{\varphi_{1}}\cup\mathbb{T}_{\varphi_{2}}$.
  • If $\varphi^{\prime}=\mathrm{F}_{[a,b]}\varphi$, then we (i) introduce a new time variable $\lambda_{i}$; (ii) add a new time constraint $\mathbb{T}_{\varphi^{\prime}}=\mathbb{T}_{\varphi}\cup\{\lambda_{i}\!\in\![a,b]\}$; (iii) increase each time indices in each progress by $\lambda_{i}$, i.e., $\mathbb{P}_{\varphi^{\prime}}=\{\mathcal{P}(c_{\Lambda}+\lambda_{i},d_{\Lambda}+\lambda_{i},\mu)\mid\mathcal{P}(c_{\Lambda},d_{\Lambda},\mu)\!\in\!\mathbb{P}_{\varphi}\}$.
  • If $\varphi^{\prime}=\mathrm{G}_{[a,b]}\varphi$, then (i) the time constraints remain unchanged, i.e., $\mathbb{T}_{\varphi^{\prime}}=\mathbb{T}_{\varphi}$; (ii) modify each invariance progress $\mathcal{I}(c_{\Lambda},d_{\Lambda},\mu)\in\mathbb{P}_{\varphi}$ to $\mathcal{I}(c_{\Lambda}+a,d_{\Lambda}+b,\mu)$ in $\mathbb{P}_{\varphi^{\prime}}$; (iii) modify each reachability progress $\mathcal{R}(c_{\Lambda},d_{\Lambda},\mu)\in\mathbb{P}_{\varphi}$, to $b-a$ progresses $\mathcal{R}(c_{\Lambda}+k,d_{\Lambda}+k,\mu)$ in $\mathbb{P}_{\varphi^{\prime}}$, where $k=a,a+1,\dots,b$.
  • If $\varphi^{\prime}=\phi\mathrm{U}_{[a,b]}\varphi$, then (i) introduce a new time variable $\lambda_{i}$; and (ii) add a new time constraint, i.e., $\mathbb{T}_{\varphi^{\prime}}=\mathbb{T}_{\varphi}\cup\{\lambda_{i}\in[a,b]\}$; and (iii) increase each time indices in each progress in $\mathbb{P}_{\varphi}$ by $\lambda_{i}$ and modify each invariance progress $\mathcal{I}(c_{\Lambda},d_{\Lambda},\mu)\in\mathbb{P}_{\phi}$ to $\mathcal{I}(c_{\Lambda}+a,d_{\Lambda}+\lambda_{i},\mu)$, i.e., $\mathbb{P}_{\varphi^{\prime}}=\{\mathcal{P}(c_{\Lambda}+\lambda_{i},d_{\Lambda}+\lambda_{i},\mu)\mid\mathcal{P}(c_{\Lambda},d_{\Lambda},\mu)\in\mathbb{P}_{\varphi}\}\cup\{\mathcal{I}(c_{\Lambda}+a,d_{\Lambda}+\lambda_{i},\mu)\mid\mathcal{I}(c_{\Lambda},d_{\Lambda},\mu)\in\mathbb{P}_{\phi}\}$.
To illustrate the above progress decomposition process, we consider the following STL formulae
$\varphi=\mathrm{F}_{[5,12]}(\mathrm{F}_{[7,16]}\mu_{1}\wedge\mathrm{G}_{[2,10]}\mu_{2})\wedge\mathrm{G}_{[18,20]}\mathrm{F}_{[4,10]}\mu_{3}.$ (9)
The decomposition process is shown in Figure 2, where the progress and time constraints are constructed incrementally from the bottom to the top. The top node represents the overall decomposed $(\mathbb{P}_{\varphi},\mathbb{T}_{\varphi})$ for the STL formula $\varphi$.
Figure 2: Decomposition Process of STL Formula (9)
@@ 3.3 Progress Allocation
Without considering the system dynamics, the STL task planning problem is essentially a constraint satisfaction problem for the decomposed progresses and constraints $(\mathbb{P}_{\varphi},\mathbb{T}_{\varphi})$. However, the unknown system dynamics introduce additional challenges, as allocating progress arbitrarily may not be feasible for the system. To address this issue, we use a search algorithm to allocate the progresses, where the feasibility of each assignment is determined based on the trajectory data.
To perform the search-based allocation process, we further split the progresses $\mathbb{P}_{\varphi}$ as follows. For each invariance progress $\mathcal{I}(a_{\Lambda},b_{\Lambda},\mu)\in\mathbb{P}_{\varphi}$, we decompose it into a reachability progress $\mathcal{R}(a_{\Lambda},a_{\Lambda},\mu)$ and an invariance progress $\mathcal{I}(a_{\Lambda}+1,b_{\Lambda},\mu)$. For simplicity, we will denote the further decomposed progresses as $(\mathbb{P},\mathbb{T})$ without subscripts, where $\mathbb{P}=\mathbb{P}^{\mathcal{R}}\dot{\cup}\mathbb{P}^{\mathcal{I}}$. Due to this further decomposition, each invariance progress follows a unique reachability progress.
Main Allocation Algorithm
Algorithm 1 Main-Allocation
0:  Initial state $\mathbf{x}_{0}$, start time $t_{0}$, reachability progresses $\mathbb{P}^{\mathcal{R}}$, invariance progresses $\mathbb{P}^{\mathcal{I}}$, time variable constraints $\mathbb{T}$
0:  A valid waypoints sequence $\bar{\mathbf{s}}$ or None if no solution is found
@@
2:  current state $\mathbf{x}\leftarrow\mathbf{x}_{0}$; current time $t\leftarrow t_{0}$
3:  task sequence $\bar{\mathbf{s}}\leftarrow[(\mathbf{x},t)]$
4:  stack $\leftarrow[(\mathbf{x},t,\mathbb{P}^{\mathcal{R}},\mathbb{T},\bar{\mathbf{s}})]$
@@
6:   $(\mathbf{x},t,\mathbb{P}^{\mathcal{R}},\mathbb{T},\bar{\mathbf{s}})\leftarrow$ pop(stack)
@@
8:    return $\bar{\mathbf{s}},\mathbb{T}$  // All reachability progresses satisfied
@@
12:     $\bar{\mathbf{s}}^{\prime}\leftarrow\bar{\mathbf{s}}.(\mathbf{x}^{\prime},t^{\prime})$
@@
16:     push $\bigl{(}\mathbf{x}^{\prime},t^{\prime},{\mathbb{P}^{\mathcal{R}}}^{\prime},{\mathbb{T}}^{\prime},\bar{\mathbf{s}}^{\prime}\bigr{)}$ onto stack
@@
The main algorithm for progress allocation is presented in Algorithm 1. The algorithm employs a depth-first search (DFS) to sequentially assign satisfaction times and waypoint for each reachability progress in $\mathbb{P}^{\mathcal{R}}$. When the algorithm terminates, it returns a sequence of waypoints with associated timestamps of form $(x_{0},t_{0})(x_{1},t_{1})\dots(x_{n},t_{n})$ such that each waypoint corresponds to the satisfication of a reachability progress in $\mathbb{P}^{\mathcal{R}}$. During the search process, we maintain the current state $\mathbf{x}$, the current time step $t$, the set of remaining reachability progresses $\mathbb{P}^{\mathcal{R}}$, the set of all time constraints $\mathbb{T}$, and the searched sequence of waypoints with associated timestamps $\bar{\mathbf{s}}$.
@@
At each step, we select a remained reachability progress $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ from $\mathbb{P}^{\mathcal{R}}$ as the next target progress from the current state $(\mathbf{x},t)$. Specifically, to achieve progress $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$, function SampleState is used to determine a satisfaction time $t^{\prime}$ and a corresponding new state $\mathbf{x}^{\prime}$ such that $\mathbf{x}^{\prime}\models\mu$. Then the searched waypoint $(\mathbf{x}^{\prime},t^{\prime})$ is appended to $\bar{\mathbf{s}}$ and we remove progress $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ from $\mathbb{P}^{\mathcal{R}}$. Furthermore, the time constraints $\mathbb{T}$ are updated based on the assigned state $\mathbf{x}^{\prime}$ and time $t^{\prime}$ according to function UpdateConstraint. Finally, the algorithm proceeds to the next iteration. If $\mathbb{P}^{\mathcal{R}}$ becomes empty, it indicates that all reachability progresses have been successfully assigned a satisfaction time. If no feasible assignment can be made, the algorithm backtracks to explore alternative assignments.
In the above DFS, we use the following heuristic order to select reachability progress from $\mathbb{P}^{\mathcal{R}}$. Let $\mathbb{T}$ be a set of time constraints and $a_{\Lambda}$ be an objective function over time variables $\Lambda$. We denote by $a_{\Lambda,\mathbb{T}}^{\mathrm{max}}$ and $a_{\Lambda,\mathbb{T}}^{\mathrm{min}}$ the maximum and minimum values of $a_{\Lambda}$ under $\mathbb{T}$, respectively, which can be easily solved by Integer Linear Programming (ILP) techniques. Then during the search process, progress $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ with earlier potential deadlines $b_{\Lambda,\mathbb{T}}^{\mathrm{min}}$ are prioritized; and if two progresses have the same deadline, the one with the earlier potential start time $a_{\Lambda,\mathbb{T}}^{\mathrm{min}}$ is preferred.
Constraint Update
@@
Let $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ be the selected progress and $(\mathbf{x}^{\prime},t^{\prime})$ be the assigned waypoint with time. Then we added the following time constraints to $\mathbb{T}$:
$a_{\Lambda}\leq t^{\prime}\quad\text{ and }\quad t^{\prime}\leq b_{\Lambda}.$ (10)
@@
Recall that in our decomposition, each original invariance progress $\mathcal{I}(a_{\Lambda},b_{\Lambda},\mu)$ is decomposed to $\mathcal{R}(a_{\Lambda},a_{\Lambda},\mu)$ and $\mathcal{I}(a_{\Lambda}+1,b_{\Lambda},\mu)$. Therefore, if constraints (10) are added for $\mathcal{R}(a_{\Lambda},a_{\Lambda},\mu)$, it means that the value of $a_{\Lambda}$ is determined as $t^{\prime}$, and we denoted by $\mathbb{P}^{\mathcal{I}}_{det}$ the set of invariance progresses whose starting times are determined. Then for those invariance progress $\mathcal{I}(c,d_{\Lambda},\mu)\in\mathbb{P}^{\mathcal{I}}_{det}$ whose starting time is determined, if $x^{\prime}\nvDash\mu$, then we further add a new constraint $d_{\Lambda}<t^{\prime}$, i.e., the invariance progress can only be effective before $t^{\prime}$ to avoid conflict with the assigned waypoint.
Sample Timed Waypoints
@@
5:     for all $\mathcal{I}(c,d_{\Lambda},\mu)\in\mathbb{P}^{\mathcal{I}}_{det}$ with determined starting time do
6:        if $\mathbf{x}^{\prime}\nvDash\mu$ then
@@
11:     if $t^{\prime}>t_{\mathrm{max}}$ or $[\max\{t^{\prime},t_{\mathrm{min}}\},t_{\mathrm{max}}]\setminus\mathcal{O}=\emptyset$ then
@@
14:     $t_{\mathrm{new}}\leftarrow$earliest time in $[\max\{t^{\prime},t_{\mathrm{min}}\},t_{\mathrm{max}}]\setminus\mathcal{O}$
@@
When a reachability progress $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ is selected, we use function SampleState to determine a valid satisfaction time and waypoint state for this progress while ensuring compliance with invariance progresses. The pseudocode of this fucntion is shown in Algorithm 2. The process starts by computing the largest possible time interval $[t_{\mathrm{min}},t_{\mathrm{max}}]$ for the reachability progress. Then the algorithm attempts to sample a candidate state $\mathbf{x}^{\prime}$ with the satisfication region of $\mu$ up to $N_{\mathrm{max}}$ times. For each attempt, we first calculate the minimum possible conflict time interval, denoted by $\mathcal{O}$, during which $\mathbf{x}^{\prime}$ conflicts with the invariance progresses. Here, we only consider invariance progresses whose starting times are determined. This is because, in the STL decomposition, each invariance progress follows a “preceding” reachability progress. If the starting time of an invariance progress is not determined, then it implies that its “preceding” reachability progress has not yet been satisfied and will only be satisfied strictly later than the current reachability progress. Consequently, this invariance progress will also start strictly later.
@@
Once the conflict time interval $\mathcal{O}$ is computed, we further use function TimePredict to predict the arrival time $t^{\prime}$ from the current state $\mathbf{x}$ to the sample state $\mathbf{x}^{\prime}$. Particularly, if (i) $t^{\prime}>t_{\mathrm{max}}$; or (ii) the feasible interval $[\max(t^{\prime},t_{\mathrm{min}}),t_{\mathrm{max}}]$ is fully occupied by conflicting intervals in $\mathcal{O}$, then it means that the sample state $\mathbf{x}^{\prime}$ is not feasible and we proceed to the next attempt. Otherwise, the earliest available time $t_{\mathrm{new}}$ in the feasible interval is assigned as the satisfaction time, and the algorithm returns $t_{\mathrm{new}}$ along with the sampled state $\mathbf{x}^{\prime}$.
Prediction of Reachability Time

### v2

@@ 3.2 Decompositions of STL Formulae
  • Invariance progress $\mathcal{I}(a_{\Lambda}+t,b_{\Lambda}+t,\mu)$: for all $t^{\prime}\in[a_{\Lambda}+t,b_{\Lambda}+t]$ it holds that $\mathbf{x}_{t^{\prime}}\vDash\mu$.
Unless stated otherwise we take $t=0$ and write simply $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ and $\mathcal{I}(a_{\Lambda},b_{\Lambda},\mu)$.
Time Variables and Constraints. The subscript $\Lambda$” indicates that the interval endpoints may depend on a finite set of time variables $\Lambda_{\varphi}=\{\lambda_{1},\ldots,\lambda_{|\Lambda|}\}$. A concrete assignment is a vector $\boldsymbol{\lambda}=[\lambda_{1},\ldots,\lambda_{|\Lambda|}]\in\mathbb{Z}_{+}^{|\Lambda|}$. Throughout our decomposition (see Appendix D), each endpoint is an affine 0-11 combination of these variables (plus constants), hence we view $a_{\Lambda},\ b_{\Lambda}:\ \mathbb{Z}_{+}^{|\Lambda|}\to\mathbb{Z}_{+}.$ All time constraints are unary interval bounds of the form $\lambda_{i}\in[\underline{a}_{i},\overline{b}_{i}]$; we collect them in a set $\mathbb{T}_{\varphi}$.
Feasible Assignment Set. The feasible set induced by $\mathbb{T}_{\varphi}$ is
$\mathcal{F}_{\varphi}:=\Big\{\boldsymbol{\lambda}\in\mathbb{Z}_{+}^{|\Lambda_{\varphi}|}\ \big|\ \forall\mathfrak{t}\in\mathbb{T}_{\varphi}:\ \boldsymbol{\lambda}\vDash\mathfrak{t}\Big\},$
where $\boldsymbol{\lambda}\vDash(\lambda_{i}\in[\underline{a}_{i},\overline{b}_{i}])$ iff the $i$-th component of $\boldsymbol{\lambda}$ lies in the specified interval. Because constraints are unary, $\mathcal{F}_{\varphi}$ factorizes coordinatewise.
Satisfaction of Progresses under an Assignment. Given $\boldsymbol{\lambda}\in\mathcal{F}_{\varphi}$, we abbreviate $\mathbf{s}_{0}\vDash\mathcal{P}(\boldsymbol{\lambda})$ for the satisfaction of $\mathcal{P}(a_{\Lambda}(\boldsymbol{\lambda}),b_{\Lambda}(\boldsymbol{\lambda}),\mu)$, where $\mathcal{P}\in\{\mathcal{R},\mathcal{I}\}$:
$\mathbf{s}_{0}\vDash\mathcal{R}(a_{\Lambda}(\boldsymbol{\lambda}),b_{\Lambda}(\boldsymbol{\lambda}),\mu)$ $\ \Leftrightarrow\ \exists t\in[a_{\Lambda}(\boldsymbol{\lambda}),b_{\Lambda}(\boldsymbol{\lambda})]:\ \mathbf{x}_{t}\vDash\mu,$ (3)
$\mathbf{s}_{0}\vDash\mathcal{I}(a_{\Lambda}(\boldsymbol{\lambda}),b_{\Lambda}(\boldsymbol{\lambda}),\mu)$ $\ \Leftrightarrow\ \forall t\in[a_{\Lambda}(\boldsymbol{\lambda}),b_{\Lambda}(\boldsymbol{\lambda})]:\ \mathbf{x}_{t}\vDash\mu.$ (4)
The full recursive decomposition procedure is detailed in Appendix D. To establish the soundness of the decomposition process with respect to the original STL specification, we propose the following lemma and the proof is provided in Appendix G.
Lemma 1 (Soundness of Semantics-based STL Decomposition).
Let $\varphi$ be an STL formula in positive normal form (PNF) without disjunctions, and let $(\mathbb{P}_{\varphi},\mathbb{T}_{\varphi})$ along with the time-variable set $\Lambda_{\varphi}$ be the result of the recursive decomposition. Let $\mathcal{F}_{\varphi}$ denote the set of feasible assignments that satisfy all time constraints. Then, for any signal $\mathbf{s}_{0}$,
$\mathbf{s}_{0}\vDash\varphi\;\Longleftrightarrow\;\exists\,\boldsymbol{\lambda}\in\mathcal{F}_{\varphi}\;\text{s.t.}\;\forall\mathcal{P}\in\mathbb{P}_{\varphi}:\;\mathbf{s}_{0}\vDash\mathcal{P}(\boldsymbol{\lambda}).$
In particular, if $\Lambda_{\varphi}=\varnothing$, i.e., no time variables appear in the decomposition procedure, the condition reduces to $\mathbf{s}_{0}\vDash\varphi\;\Longleftrightarrow\;\forall\mathcal{P}\in\mathbb{P}_{\varphi}:\;\mathbf{s}_{0}\vDash\mathcal{P}.$
@@ 3.3 Progress Allocation
Basic Idea. By Lemma 1, when system dynamics are ignored, STL planning reduces to a constraint-satisfaction problem over the decomposed progresses and time constraints $(\mathbb{P}_{\varphi},\mathbb{T}_{\varphi})$. Concretely, $\mathbf{s}_{0}\vDash\varphi$ holds iff there exists a feasible time-variable assignment $\boldsymbol{\lambda}\in\mathcal{F}_{\varphi}$ such that every progress in $\mathbb{P}_{\varphi}$ is satisfied. This reformulation is advantageous for signal construction: instead of reasoning directly with temporal quantifiers in the original formula, it suffices to (i) ensure the existence of a feasible $\boldsymbol{\lambda}$, and (ii) construct a signal that satisfies all progresses under that assignment. We refer to this process as progress allocation.
However, unknown system dynamics pose additional challenges, as arbitrary allocation may result in dynamically infeasible plans. To address this, we adopt a search-based allocation algorithm, where the feasibility of each assignment is evaluated using a model (Time Predictor) learned from trajectory data, which implicitly captures the system dynamics.
Preprocessing. To perform the search-based allocation process, we further split the progresses $\mathbb{P}_{\varphi}$ as follows. For each invariance progress $\mathcal{I}(a_{\Lambda},b_{\Lambda},\mu)\in\mathbb{P}_{\varphi}$, we decompose it into a reachability progress $\mathcal{R}(a_{\Lambda},a_{\Lambda},\mu)$ and an invariance progress $\mathcal{I}(a_{\Lambda}+1,b_{\Lambda},\mu)$. For simplicity, we will denote the further decomposed progresses as $(\mathbb{P},\mathbb{T})$ without subscripts, where $\mathbb{P}=\mathbb{P}^{\mathcal{R}}\dot{\cup}\mathbb{P}$$^{\mathcal{I}}$. Due to this further decomposition, each invariance progress follows a unique reachability progress.
Main Allocation Algorithm. The main algorithm for progress allocation is presented in Algorithm 1. Note that the pseudo-code presents the search with an explicit stack for readability, whereas our reference implementation executes the same logic recursively.
Algorithm 1 Main‑Allocation
0:  Initial state $\mathbf{x}_{0}$, reachability progresses $\mathbb{P}^{\mathcal{R}}$, invariance progresses $\mathbb{P}^{\mathcal{I}}$, time variable constraints $\mathbb{T}$
0:  A valid waypoints sequence $\tilde{\mathbf{s}}$ or None if no solution is found
@@
2:  current state $\mathbf{x}\leftarrow\mathbf{x}_{0}$; current time $t\leftarrow 0$
3:  sequence $\tilde{\mathbf{s}}\leftarrow[(\mathbf{x},t)]$
[…]
@@
[…]
@@
8:        return $\tilde{\mathbf{s}},\mathbb{T}$    // All reachability progresses satisfied
@@
12:           $\tilde{\mathbf{s}}^{\prime}\leftarrow\tilde{\mathbf{s}}.(\mathbf{x}^{\prime},t^{\prime})$
@@
[…]
@@
The algorithm employs a depth-first search (DFS) to sequentially assign satisfaction times and waypoints for each reachability progress in $\mathbb{P}^{\mathcal{R}}$. When the algorithm terminates, it returns a sequence of waypoints with associated timestamps of form $(\tilde{\mathbf{x}}_{0},t_{0})(\tilde{\mathbf{x}}_{1},t_{1})\dots(\tilde{\mathbf{x}}_{n},t_{n})$ such that each waypoint corresponds to the satisfaction of a reachability progress in $\mathbb{P}^{\mathcal{R}}$. During the search process, we maintain the current state $\mathbf{x}$, the current time step $t$, the set of remaining reachability progresses $\mathbb{P}^{\mathcal{R}}$, the set of all time constraints $\mathbb{T}$, and the searched sequence of waypoints with associated timestamps $\tilde{\mathbf{s}}$.
@@
At each step, we select a remained reachability progress $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ from $\mathbb{P}^{\mathcal{R}}$ as the next target progress from the current state $(\mathbf{x},t)$. Specifically, to achieve progress $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$, function SampleState is used to determine a satisfaction time $t^{\prime}$ and a corresponding new state $\mathbf{x}^{\prime}$ such that $\mathbf{x}^{\prime}\models\mu$. Then the searched waypoint $(\mathbf{x}^{\prime},t^{\prime})$ is appended to $\tilde{\mathbf{s}}$ and we remove progress $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ from $\mathbb{P}^{\mathcal{R}}$. Furthermore, the time constraints $\mathbb{T}$ are updated based on the assigned state $\mathbf{x}^{\prime}$ and time $t^{\prime}$ according to function UpdateConstraint. Finally, the algorithm proceeds to the next iteration. If $\mathbb{P}^{\mathcal{R}}$ becomes empty, it indicates that all reachability progresses have been successfully assigned a satisfaction time. If no feasible assignment can be made, the algorithm backtracks to explore alternative assignments.
Dynamic Maintenance of Time Variable Constraints. During allocation, we dynamically maintain the set of time variable constraints $\mathbb{T}$. Once the satisfaction time of a reachability progress is determined, additional constraints are added accordingly. This enables the planner to query $\mathbb{T}$ at any point to determine potential ranges of $a_{\Lambda}$ and $b_{\Lambda}$ (we denote by $a_{\Lambda,\mathbb{T}}^{\mathrm{min}},a_{\Lambda,\mathbb{T}}^{\mathrm{max}}],[b_{\Lambda,\mathbb{T}}^{\mathrm{min}},b_{\Lambda,\mathbb{T}}^{\mathrm{max}}]$) in a progress $\mathcal{P}(a_{\Lambda},b_{\Lambda},\mu)$ by solving integer linear programs (ILP), while ensuring consistency with previously assigned progresses. For example, to obtain $a_{\Lambda,\mathbb{T}}^{\mathrm{min}}$, we minimize $a_{\Lambda}(\boldsymbol{\lambda})$ subject to all constraints in $\mathbb{T}$.
Note that only the constraints are updated during planning; the time variable set $\Lambda$ remains fixed. As new constraints are introduced, the feasible assignment set $\mathcal{F}$ becomes increasingly restricted. If $\mathcal{F}$ becomes empty-i.e., no assignment $\boldsymbol{\lambda}$ satisfies all constraints-the current allocation is deemed infeasible, and backtracking is triggered to revise earlier decisions.
Heuristic Order. In the DFS of Algorithm 1, reachability progresses from $\mathbb{P}^{\mathcal{R}}$ are selected based on a heuristic order: progresses with earlier potential deadlines $b_{\Lambda,\mathbb{T}}^{\mathrm{min}}$ are prioritized; if multiple candidates share the same deadline, the one with the earlier potential start time $a_{\Lambda,\mathbb{T}}^{\mathrm{min}}$ is preferred.
@@
Constraint Update. Let $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ be the selected progress and $(\mathbf{x}^{\prime},t^{\prime})$ the assigned waypoint with time. We then add the following time constraints to $\mathbb{T}$:
$a_{\Lambda}\leq t^{\prime}\quad\text{and}\quad b_{\Lambda}\geq t^{\prime}.$ (5)
@@
These constraints ensure that the progress can be satisfied at time $t^{\prime}$. Recall that in our decomposition, each original invariance progress $\mathcal{I}(a_{\Lambda},b_{\Lambda},\mu)$ is split into a reachability progress $\mathcal{R}(a_{\Lambda},a_{\Lambda},\mu)$ and a residual invariance progress $\mathcal{I}(a_{\Lambda}+1,b_{\Lambda},\mu)$. Thus, if constraints (5) are added for a reachability progress of the form $\mathcal{R}(a_{\Lambda},a_{\Lambda},\mu)$, the value of $a_{\Lambda}$ is fixed to $t^{\prime}$. Let $\mathbb{P}^{\mathcal{I}}_{\text{det}}$ denote the set of invariance progresses whose starting times have been determined. For any $\mathcal{I}(c,d_{\Lambda},\mu)\in\mathbb{P}^{\mathcal{I}}_{\text{det}}$, if $\mathbf{x}^{\prime}\nvDash\mu$, we add an additional constraint $d_{\Lambda}<t^{\prime}$ to ensure that the invariance progress terminates before $t^{\prime}$, thereby avoiding conflicts with the assigned waypoint.
Sample Timed Waypoints. When a reachability progress $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ is selected, we use function SampleState to determine a valid satisfaction time and waypoint state for this progress while ensuring compliance with invariance progresses.
@@
5:   for all $\mathcal{I}(c,d_{\Lambda},\mu^{\prime})\in\mathbb{P}^{\mathcal{I}}_{det}$ with determined starting time do
6:    if $\mathbf{x}^{\prime}\nvDash\mu^{\prime}$ then
@@
11:   if $t^{\prime}>t_{\mathrm{max}}$ or $\max\{t^{\prime},t_{\mathrm{min}}\},t_{\mathrm{max}}]\setminus\mathcal{O}=\emptyset$ then
@@
14:   $t_{\mathrm{new}}\leftarrowearliest$ time in $\max\{t^{\prime},t_{\mathrm{min}}\},t_{\mathrm{max}}]\setminus\mathcal{O}$
@@
The pseudocode of this function is shown in Algorithm 2. The process starts by computing the largest possible time interval $t_{\mathrm{min}},t_{\mathrm{max}}]$ for the reachability progress. Then the algorithm attempts to sample a candidate state $\mathbf{x}^{\prime}$ with the satisfication region of $\mu$ up to $N_{\mathrm{max}}$ times. For each attempt, we first calculate the minimum possible conflict time interval, denoted by $\mathcal{O}$, during which $\mathbf{x}^{\prime}$ conflicts with the invariance progresses.
@@
Once the conflict time interval $\mathcal{O}$ is computed, we further use function TimePredict to predict the arrival time $t^{\prime}$ from the current state $\mathbf{x}$ to the sample state $\mathbf{x}^{\prime}$. Particularly, if (i) $t^{\prime}>t_{\mathrm{max}}$; or (ii) the feasible interval $\max(t^{\prime},t_{\mathrm{min}}),t_{\mathrm{max}}]$ is fully occupied by conflicting intervals in $\mathcal{O}$, then it means that the sample state $\mathbf{x}^{\prime}$ is not feasible and we proceed to the next attempt. Otherwise, the earliest available time $t_{\mathrm{new}}$ in the feasible interval is assigned as the satisfaction time, and the algorithm returns $t_{\mathrm{new}}$ along with the sampled state $\mathbf{x}^{\prime}$.
Remark 1.
When calculating the conflict time interval $\mathcal{O}$, we only consider invariance progresses whose starting times are determined. This is because, in the preprocessing process, we make each invariance progress follows a “preceding" reachability progress. If the starting time of an invariance progress is not determined, then it implies that its “preceding" reachability progress has not yet been satisfied and will only be satisfied strictly later than the current reachability progress. Consequently, this invariance progress will also start strictly later.